### v1

# Wave pulses with unusual asymptotical behavior at infinity

Abstract
The behavior of wave signals in the far zone is not only of theoretical interest but also of paramount practical importance in communications and other fields of applications of optical, electromagnetic or acoustic waves. Long time ago T. T. Wu introduced models of “electromagnetic missiles” whose decay could be made arbitrarily slower than the usual inverse distance by an appropriate choice of the high frequency portion of the source spectrum. Very recent work by Plachenov and Kiselev introduced a finite-energy scalar wave solution, different from Wu’s, decaying slower than inversely proportional with the distance. A physical explanation for the unusual asymptotic behavior of the latter will be given in this article. Furthermore, two additional examples of scalar wave pulses characterized by abnormal slow decay in the far zone will be given and their asymptotic behavior will be discussed. A proof of feasibility of acoustic and electromagnetic fields with the abnormal asymptotics will be described.
wave equation, unidirectional wavepackets, ultrashort electromagnetic and acoustical pulses
preprint: APS/123-QED
I Introduction
It is a textbook truth that solutions of the wave equation decay inversely proportional to distance in the far field or the wave zone (see, e.g., [1, 2]). However, a long time ago Wu [3] showed that the decay can be made arbitrarily slower by a careful choice of the high-frequency wing of the source spectrum. In the 1980s, closely related to Wu’s ”electromagnetic missiles”, another subject dealing with uncommon solutions of the wave equation emerged—the so-called localized or non-diffracting or space-time wave packets possessing exotic properties, such as self-healing, superluminality, and invariant propagation without spread or decay, theoretically up to infinity. By now this subject has grown into a large research field of its own, see monograph [4] and reviews [5, 6, 7, 8, 9]. Therefore, the decay in the wave zone as a specific subject in mathematical physics has remained somewhat in the shadows. However, in recent years, several papers on the study of asymptotical behavior have appeared, see [10, 11, 12] and Refs. therein.
The very recent paper by Plachenov and Kiselev [12] introduced a finite-energy solution, different from Wu’s model [3], that decays slower than inversely proportional to the distance.
They consider the following simple solution of the homogeneous wave equation (we have somewhat changed the designations)
$\psi\left(R,t\right)=\frac{1}{\left(ct+ict_{s}\right)^{2}-R^{2}},$ (1)
where $R=\left(x^{2}+y^{2}+z^{2}\right)^{1/2}$ is the distance of the field point from the origin, $c$ is the velocity of light (or sound), and $t_{s}$ is a constant that determines the width of the pulse. $\psi\left(R,t\right)$ is a special case of a splash pulse [13]. It is frequently used for modeling ultrashort pulses and constitutes a spherically propagating pulse, while the spherical surface, where the pulse peaks at $R\approx ct,$ collapses into the origin at negative times $t<0$ and expands for $t>0.$ The real part $\operatorname{Re}\psi\left(R,t\right)$ represents a bipolar pulse and the imaginary part $\operatorname{Im}\psi\left(R,t\right)$ a unipolar one everywhere where $R\gg ct_{s}$, and vice versa near the origin where $R<ct_{s}$.
They consider the limit
$L=\lim_{t\rightarrow\infty}\left[ct\,\psi\left(R,t\right)\right]_{R=ct+\Delta}\ ,$ (2)
where $\Delta$ is zero or an arbitrary (not large) real number, so that substitution $R=ct+\Delta$ leads to so to speak riding on the top of the pulse peak or its proximity as $t\rightarrow\infty$. For a wave function in the form of Eq. (1) it turns out that $L=1/\left(2ict_{s}-2\Delta\right)$, i.e., it is finite.
The limit of the type of Eq. (2) plays an important role in various fields of mathematical physics and it is known that for finite-energy ”well-behaving” solutions of the wave equation the limit $L$ should be finite (incl. $L=0$), see, e.g., [1, 2, 10, 11].
The authors of [12] showed that for the primitive (antiderivative) of $\psi\left(R,t\right)$
[…]
in contradistinction to the usual behaviour of fields in the wave zone, the quantity in Eq. (2) is diverging:
$L=-\left(1/2c\right)\lim_{t\rightarrow\infty}\ln ct=-\infty.$ (4)
They also showed that $\Psi\left(R,t\right)$ has finite total energy, i.e., it is not an unphysical solution.
The purpose of our article is to present additional examples of pulses with unusual asymptotics and to reveal physical causes of abnormally slow decay. In the next section we consider one of such examples—the primitive (antiderivative) of the wave function of a unidirectional pulse. Sect. III is devoted to another example—the fractional splash pulse which already itself, without taking its time integral, decays very slowly in the far zone.
The authors in the short communication [12] neither illustrate graphically the behavior of $\Psi\left(R,t\right)$ nor interpret physically its unusual asymptotics. In Sec. IV we try to fill in this gap. In Sect. V we discuss some subtleties of the slow decay of the other pulses studied in Sect. II and III and consider the physical feasibility in optics, electromagnetics, or acoustics of the pulses considered in this study.
II Hemispherical unidirectional pulse
Now we will consider a cylindrically symmetric pulse which is unidirectional in the sense that all its plane-wave constituents propagate solely into the hemisphere with the axial component of the wave vector $k_{z}>0$. Such pulses have been intensively studied in recent years [14, 15, 16, 17, 18, 19]. Surprisingly, a unidirectional pulse may exhibit locally negative values of the $z$ component of the Poynting vector, i.e., energy flow in the direction opposite to the pulse propagation axis $z$. Such an energy backflow effect was studied in detail with respect to a class of pulses a typical representative of which is given by the following expression [17]:
$\begin{split}u\left(\rho,z,t\right)&=\frac{1}{g\left(\rho,t\right)\left[i\left(z-iz_{s}\right)+g\left(\rho,t\right)\right]},\\ g\left(\rho,t\right)&\equiv\sqrt{\rho^{2}-\left(ct+ict_{s}\right)^{2}}.\end{split}$ (5)
Here $z_{s}$ is a positive constant that contributes additionally to the pulsewidth. It is interesting to note that $u\left(\rho,z,t\right)$ is a double antiderivative, first with respect to time and then with respect to the variable $z$, of another unidirectional solution, given in Eq. (5) of Ref. [17] (with replacement $z\rightarrow-z$), which in turn is a compact version of Eq. (3.10) derived in Ref. [14] by a unidirectional Fourier synthesis.
Similarly to $\psi\left(R,t\right)$ in Eq. (1), the real and imaginary parts of $u\left(R,t\right)$ represent bipolar and unipolar pulses, but in contrast to the spherically symmetric case, in the given case the pulses ”ride” on a collapsing-expanding tube-like surface with a hemispherical surface inside (see plots below).
$u\left(\rho,z,t\right)$ has normal asymptotics:
$\lim_{t\rightarrow\infty}\left[ct\,u\left(\rho,z,t\right)\right]_{z=ct+\Delta}=\frac{i}{ct_{s}+z_{s}+i\Delta},$ (6)
$\lim_{t\rightarrow\infty}\left[ct\,u\left(\rho,z,t\right)\right]_{z=-(ct+\Delta)}=0,$ (7)
$\lim_{t\rightarrow\infty}\left[ct\,u\left(\rho,z,t\right)\right]_{\rho=ct+\Delta}=\frac{i}{2ct_{s}+2i\Delta}.$ (8)
But how does its primitive behave? Taking the integral of $u\left(\rho,z,t\right)$ is a complicated task. It can be found neither in tables of integrals nor can be taken by the newest versions of computation packages Mathematica and Maple. The integral over time from $\infty$- ∞ to $t$ was taken first by setting $t_{s}=z_{s}=0$ and then making replacements $z\rightarrow z-iz_{s}$ and $t\rightarrow t+it_{s}$. After some algebra, we found the following expression for the primitive, where $g\left(\rho,t\right)$ is given in Eq. (5).
$U\left(\rho,z,t\right)$ $=\frac{1}{2h\left(\rho,z\right)}\left[U_{1}\left(\rho,z,t\right)+U_{2}\left(\rho,z,t\right)\right],$ (9)
$U_{1}\left(\rho,z,t\right)$ $\equiv\ln\frac{ct+ict_{s}+h\left(\rho,z\right)}{ct+ict_{s}-h\left(\rho,z\right)},\qquad h\left(\rho,z\right)\equiv\sqrt{\rho^{2}+\left(z-iz_{s}\right)^{2}},$
$U_{2}\left(\rho,z,t\right)$ $\equiv\ln\frac{\left[\left(-ict+ct_{s}\right)\left(z-iz_{s}\right)+g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z-iz_{s}-h\left(\rho,z\right)\right]}{\left[\left(-ict+ct_{s}\right)\left(z-iz_{s}\right)-g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z-iz_{s}+h\left(\rho,z\right)\right]}.$
The behavior of $U\left(\rho,z,t\right)$ is depicted in Figs. 1 and 2. We see that the pulse peaks on the surface of a collapsing-expanding tube with a hemisphere inside it, while the radius of both is equal to $|ct|$. Hence, $U\left(\rho,z,t\right)$ behaves similarly to $u\left(\rho,z,t\right)$ from Eq. (5), the most remarkable difference being that the peak of the real part of the latter is abruptly and symmetrically bipolar, except at $t=0$ when it is unipolar contrary to Fig. 1.
$U\left(\rho,z,t\right)$ has like $\Psi\left(R,t\right)$ abnormal asymptotics, except when the limiting process runs in the negative direction of the axis $z$:
$\lim_{t\rightarrow\infty}\left[ct\,U\left(\rho,z,t\right)\right]_{z=ct+\Delta}=\infty~{},$ (10)
$\lim_{t\rightarrow\infty}\left[ct\,U\left(\rho,z,t\right)\right]_{z=-(ct+\Delta)}=\ln 2~{},$ (11)
$\lim_{t\rightarrow\infty}\left[ct\,U\left(\rho,z,t\right)\right]_{\rho=ct+\Delta}=\infty~{},$ (12)
Another exception where the limit proves to be finite and equals to $\ln 2$ is when time $t$ runs to $\infty$- ∞ but $z$ to +$\infty$+ ∞. In both these exceptional cases the pulse peak is not involved into the process, as can be seen by looking at Figs. 1 and 2.
We calculated the Poynting vector and energy density of $U\left(\rho,z,t\right)$ according to known formulas (see, e.g., Eqs. (1)-(5) of Ref. [18]) for scalar fields, as well as the total energy. The latter is finite. The former two have normal asymptotics in the sense of finite values of the limit of the type in Eq. (2) as they have to be since both consist of derivatives of $U\left(\rho,z,t\right)$. Despite the unidirectionality of $U\left(\rho,z,t\right)$—integration as a linear operation preserves the unidirectional spectrum of $u\left(\rho,z,t\right)$— in certain spatial regions the energy flux turned out to be directed opposite to the azis $z$, and at some points the velocity of such an energy backflow even reached almost the value $-c$.
We calculated the electric and magnetic fields from $U\left(\rho,z,t\right)$ using the Hertz vector technique and proved that they have normal asymptotics. This is understandable, as the technique involves taking derivatives of the scalar field.
Finally, we found that the imaginary part of $U\left(\rho,z,t\right)$ is a ”strange” field [20, 21, 22] in the sense that in certain regions the time integral from $\infty$- ∞ to +$\infty$+ ∞ is not zero. This is understandable, because the imaginary part is a unipolar pulse.
III Fractional splash pulse
Members of a general family of splash pulses are derived as spectral superpositions of the most investigated localized propagation-invariant unipolar pulse—the focus wave mode [5, 6, 4]. A specific fractional splash pulse is given by the following expression
$f\left(\rho,z,t\right)$ $=\frac{1}{a_{1}+i\left(z-ct\right)}$ (13)
×$\times\frac{1}{\left(a_{2}-i\left(z+ct\right)+\rho^{2}\left(a_{1}+i\left(z-ct\right)\right)^{-1}\right)^{3/4}},$
where $a_{1}$ and $a_{2}$ are positive parameters. The general form of the splash mode pulse is given by the expression in Eq. (13), with the power $3/4$ in the denominator replaced by $\nu+1$, with $\nu>-1$. The fractional pulse above corresponds to $\nu=-1/4$.
The behavior of $f\left(\rho,z,t\right)$ is depicted in Fig. 3. Both the real and imaginary parts of the pulse possess a collapsing-expanding spherical structure, but a peak appears only in a narrow cone of one hemisphere (of positive $z$ if $ct>0$ and of negative $z$ if $ct<0$). The real part has mirror symmetry with respect to time reversal $t\rightarrow-t$, while the imaginary part additionally reverses its sign. Outside the region of the origin, the pulse tops of the real part and the imaginary part do not coincide: the top of the former is shifted farther from the origin $z=0$ than the point $z=ct$ by a generally small distance about $32a_{1}/$ 3, and the top of the imaginary part is shifted closer by the distance about $a_{1}/4$. This follows from a detailed inspection of Eq. (13), as well as from plots which will be presented in Section V. But the most remarkable distinction from the pulses considered in the previous sections is that the fractional splash pulse itself, without the need to take its antiderivative, exhibits abnormal asymptotics as the limits of the type in Eq. (2) show. The limits along different directions turn out to be as follows.
$\lim_{t\rightarrow\infty}\left[ct\,f\left(\rho,z,t\right)\right]_{z=ct+\Delta}=\infty\cdot\frac{\exp{(i\frac{3\pi}{8})}}{a_{1}+i\Delta}~{},$ (14)
$\lim_{t\rightarrow\infty}\left[ct\,f\left(\rho,z,t\right)\right]_{z=-ct-\Delta}=\frac{i}{2(a_{2}+i\Delta)^{3/4}}~{},$ (15)
$\lim_{t\rightarrow\infty}\left[ct\,f\left(\rho,z,t\right)\right]_{\rho=ct+\Delta}=\frac{i}{(a_{1}+a_{2}+2i\Delta)^{3/4}}~{},$ (16)
$\lim_{t\rightarrow\infty}\left[ct\,f\left(\rho,z,t\right)\right]_{\rho=ct+\Delta,\,z=ct+\Delta}=0~{}.$ (17)
Eq. (14) shows that the peaks of both the real and imaginary parts decay slower than the normal decay $\sim 1/z$ when one moves with the pulse top along the positive direction of the axis $z$ towards infinity. How they approach infinity will be discussed in Sect. V. As was mentioned above, the tops of the pulses of the real and imaginary parts of $f\left(\rho,z,t\right)$ are shifted from the point $z=ct$, and therefore the value of $\Delta$ affects differently the limits of the parts. If $ct>0$ and $z>0$, a simple choice is $\Delta=0$ in which case the running point is simultaneously close to the tops of the pulses of both the real and imaginary parts and the exponent in Eq. (14) shows that at this point both $ct\,\Re f\left(\rho,z=ct,t\right)$ and $ct\,\Im f\left(\rho,z=ct,t\right)$ approach infinity in the same way.
Eqs. (15)-(17) state that if $t\rightarrow+\infty$ in other direction than along the positive axis $z$, the pulse exhibits normal asymptotics. Even if one modifies Eq. (17) by replacements $\rho=ct/\sqrt{2}+\Delta,\,z=ct/\sqrt{2}+\Delta$ in order to keep the point (running along the diagonal direction) close to the pulse’s top, the expression for the limit remains finite.
We calculated the Poynting vector and energy density of $f\left(\rho,z,t\right)$. Although individually time and space derivatives of the fractional splash pulse exhibit abnormal asymptotic behavior, the energy density and the Poynting vector when inserted into Eq. (14) instead of $f$ do not result in an infinite limit because both quantities are quadratic in time and space derivatives. However, for energy-type quantities, a normal decay is inversely proportional to the square of the distance. This means that in testing the asymptotic behavior, $ct$ must be replaced by $c^{2}t^{2}$ in the expression of the limit. In doing so, we arrive at the conclusion that the energy density and the Poynting vector also decay abnormally slowly. The total energy is finite. $f\left(\rho,z,t\right)$ is square integrable, which is also important from a physical point of view as will be discussed in Sect. V.
We studied also other fractional splash pulses which are given by the same expression in Eq. (13) but with other values of the power instead of $3/4$ in the denominator. If the power is $1/2$, the pulse has the same properties as $f\left(\rho,z,t\right)$, but unfortunately its total energy is infinite. But if the value of the power is between $1/2$ and $1$, i.e., if $\frac{1}{2}<\nu+1<1$, the pulse has slower than normal decay and at the same time its wave function is square-integrable and has finite total energy.
IV Causes of uncommon asymptotical behavior at infinity
It is known that solutions to the homogeneous wave equation can be expressed by a convolution of the density of a fictitious Huygens source (coupled with sink) and the free-space propagator or the Riemann - Schwinger function
$D_{0}(t,R)$ $=G_{+}(t,R)-G_{-}(t,R)=$ (18)
$=\frac{c}{4\pi R}\left[~{}\delta\left(R-ct\right)-\delta\left(R+ct\right)~{}\right]~{},$
where $G_{+}$ and $G_{-}$ are the retarded and advanced Green functions, respectively. By doing this with the delta-point-like source function
$\rho(\textbf{r},t)=\delta(\textbf{r})\hskip 2.0pt\frac{1}{2c(t+it_{s})}~{},$ (19)
a field like $\psi\left(R,t\right)$ in Eq. (1) is expressed as a difference of converging and expanding spherical waves [23]. Specifically, $\psi\left(R,t\right)$ can be decomposed into these two waves simply by elementary algebra, viz.:
$\psi\left(R,t\right)$ $=\psi_{+}\left(R,t\right)-\psi_{-}\left(R,t\right),$ (20)
$\psi_{+}(R,t)$ $=\frac{1}{2R}\hskip 2.0pt\frac{1}{ct-R+ict_{s}}~{},$ (21)
$\psi_{-}(R,t)$ $=\frac{1}{2R}\hskip 2.0pt\frac{1}{ct+R+ict_{s}}~{}.$ (22)
Here $\psi_{-}$ represents the spherical pulse converging at negative times to the origin and $\psi_{+}$ expanding from it at positive times.
In Fig. 4 the imaginary parts of the expanding and converging waves given by Eq. (21) and Eq. (22), respectively, are plotted against the radius $R$ at three different instants of time. The curves are plotted with reverse sign for better readability and are scaled—multiplied by $R$—in order to eliminate the factor $1/R$. Thanks to the scaling we observe that the peaks of the pulses at different time instants are of equal height. This correponds to finitness of the limit in Eq. (2), and means that the splash pulse has indeed normal asymptotics, as mentioned already in Sect.1. The same holds also for the real parts of the wave functions in Eqs. (21) and (22), except that the pulse of the real part is bipolar outside of the region of origin at $t=0$, where it is unipolar, and the pulse of the imaginary part is, on the contrary, bipolar. The 50x amplified third curve in Fig. 4 shows the residual tail of the converging wave, which is very weak at $ct=30$ because its peak is already gone by positive times. We shall use Fig. 4 later as a template for graphical illustration for the explanation of the abnormal asymptotics of the primitive in Eq. (3).
Quite analogously to Eqs. (20)-(22), the primitive of the splash pulse $\Psi\left(R,t\right)$ can be decomposed into expanding and converging waves, as can be seen readily from Eq. (3)  [^1] If one integrates $\psi_{+}\left(R,t\right)$ and $\psi_{-}\left(R,t\right)$ over time from $\infty$- ∞ up to any finite value $t$, then the real parts of both $\Psi_{+}\left(R,t\right)$ and $\Psi_{-}\left(R,t\right)$ acquire an additive integration constant $C=-ln(\infty)$ which cancels out in Eq. (23) and therefore does not affect $\Psi\left(R,t\right)$., viz.:
$\Psi\left(R,t\right)$ $=\Psi_{+}\left(R,t\right)-\Psi_{-}\left(R,t\right),$ (23)
$\Psi_{+}(R,t)$ $=\frac{1}{2R}\hskip 2.0pt\ln(ct-R+ict_{s})~{},$ (24)
$\Psi_{-}(R,t)$ $=\frac{1}{2R}\hskip 2.0pt\ln(ct+R+ict_{s})~{}.$ (25)
In Fig. 5 the real parts of expanding and converging waves given, respectively, by Eq. (24) and Eq. (25), as well as of $\Psi\left(R,t\right)$ from Eq. (23), are plotted against the radius $R$ at two different instants of time. We no longer deal with the imaginary parts because their asymptotics is not abnormal. This is understandable, because as integrands in Eq. (3), the imaginary parts of $\psi_{+}\left(R,t\right)$ and $\psi_{-}\left(R,t\right)$ decay with $t\rightarrow\infty$ as $t^{-2}$, while the real parts as $t^{-1}$.
The curves are multiplied by $R$ and so become scaled appropriately for their comparison. We observe that the peaks of the real part of $\Psi_{+}(R,t)$ at different time instants are of equal height. This corresponds to finitness of the limit in Eq.(2) and means that the expanding part of the primitive, taken separately, has normal asymptotics. But we also observe, that if the tails of the converging waves are added— according to Eq. (23) with negative sign—the peaks of $R\cdot\Re\left[\Psi\left(R,t\right)\right]$ of the splash pulse grow logarithmically toward $\infty$- ∞. Indeed, the sum of the ordinate of the peak at $R=30$ of $R\cdot\Re\left[\Psi_{+}(R,30)\right]$ and that of $-R\cdot\Re\left[\Psi_{-}(R,30)\right]$ is exactly equal to the ordinate of the peak of $R\cdot\Re\left[\Psi(R,30)\right]$ at the same abscissa $R=30$. The same can be observed for the curves for the time instant $ct=160$.
The reason of the abnormal asymptotics of the primitive $\Psi(R,t)$ is the presence at $t>0$ of the tail of the converging splash pulse, which according to Eq. (3) has been integrated from $ct=-\infty$ up to a given positive time instant. Had we considered a non-fictitious radiation source switched on at $ct=0$ and used the retarded Green function only, $\Psi_{-}(R,t)$ would be left out of Eq. (23) and $\Psi(R,t)$ would have normal asymptotics.
Finally, we calculated the energy flux in the pulse $\Psi\left(R,t\right)$, i.e., the Poynting vector and found that for it the limit of type Eq. (2) is finite, as it has to be since the Poynting vector consists of derivatives of $\Psi\left(R,t\right)$.
V Discussion and study of physical feasibility
The real part of the primitive of the unidirectional pulse has according to Eq. (10) abnormal asymptotics. With $ct=z\rightarrow\infty$ the real part decays as $z^{-1}\ln[z/(ct_{s}+z_{s})]$. Unfortunately, the expression in Eq. (9) of the primitive of the unidirectional pulse cannot be decomposed into expanding and converging parts. However, it is remarkable that in the particular case $z=0$ and $z_{s}=0$ Eq. (9) simplifies substantially, viz.:
$U\left(\rho,z,t\right)\mid_{z=z_{s}=0}~{}=-\frac{1}{2c\rho}\ln\left(\frac{ct+ict_{s}-\rho}{ct+ict_{s}+\rho}\right)$ (26)
and coincides with Eq. (3) if we replace $R\rightarrow\rho$ and change the sign. Hence, all the results for the primitive of the splash pulse also apply to the radial evolution of the primitive of the unidirectional pulse.
The time dependence of the imaginary part of the fractional splash pulse $f\left(\rho,z,t\right)$ at different points on the axis $z$ is depicted in Fig. 6. We observe that in the region $ct>0$ the strength of the pulses (multiplied by $ct$) grows toward infinity, while in the region $ct<0$ it remains constant. The real part behaves analogously.
Contrary to the case of $z-$dependence with fixed time instants, in the case of $ct-$dependence with fixed locations $z$ the tops of the curves of the imaginary parts are shifted to the right (if $t>0$) from the point where $ct$ equals to a fixed value of $z$. That is why in order for the dashed curve to touch the tops of the peaks, in the expression of $ct\cdot\Im f\left(\rho=0,z=ct+\Delta,t\right)$ the value of $\Delta$ was set to $\Delta=-a_{1}/4$. If one sets $\Delta=0$, the dashed curve exactly coincides with the dotted one. The latter presents an asymptotical behavior of $ct\cdot f\left(\rho=0,z=ct+\Delta,t\right)$ according to the following expression which is directly extracted from the series expansion of the expression in Eq. (13).
$ctf\left(\rho=0,z=ct+\Delta,t\right)\propto C(a_{1},\Delta){\sqrt[4]{ct}},$
$C(a_{1},\Delta)$ $=\frac{1}{(-2i)^{3/4}(a_{1}+i\Delta)},$ (27)
$\Re\,C(a_{1},\Delta)$ $=\frac{\cos(3\pi/8)a_{1}+\sin(3\pi/8)\Delta}{2^{3/4}(a_{1}^{2}+\Delta^{2})},$ (28)
$\Im\,C(a_{1},\Delta)$ $=\frac{\sin(3\pi/8)a_{1}-\cos(3\pi/8)\Delta}{2^{3/4}(a_{1}^{2}+\Delta^{2})}.$ (29)
The dotted curve has been plotted according to Eq. (29) as $\Im\,C(a_{1},\Delta)\sqrt[4]{ct}$ with $\Delta=0$. If one puts here $\Delta=-3/4$, the dotted curve coincides exactly with the dashed curve.
It follows, then,that the the pulse propagating in the positive direction of the axis $z$ decays as $\sim 1/(z=ct)^{3/4}$, i.e., slower than the normal decay $\sim 1/(z=ct)$. As one can observe in Fig. 6, the pulse acquires the abnormal asymptotical behavior already quite close to the origin—after propagating a distance about a ten-fold of its width.
The fractional splash pulse is not unidirectional as its $k_{z}$-spectrum shows. This is understandable because it can be presented as a superposition of focus wave modes which are bidirectional (’non-causal’) pulses. The latter can be decomposed into forward and backward propagating parts, but the corresponding expressions are very complicated consisting of Lommel functions [25]. Therefore, we could not make use of the superpositional relation between the focus wave mode and the fractional splash pulse for decomposing the latter.
Not only the total energy of the scalar field $f\left(\rho,z,t\right)$ is finite, but also the total energy of an electromagnetic field which can be proved by a relation [26] between the square integrability of the spectrum of a splash pulse and the total energy.
We undertook a thorough proof of the physical feasibility of electromagnetic fields derivable from the fractional splash pulse $f\left(\rho,z,t\right)$ with the help of a Hertz and a Riemann-Silberstein vector technique. Specifically, we chose the Hertz vector as $\vec{\Pi}\left(x,y,z,t\right)=\vec{m}f\left(\rho=\sqrt{x^{2}+y^{2}},z,t\right)$ where $\vec{m}$ is a constant directional vector of type $(1,0,0)$ or $(1,1,0)$ or $(1,i,0)$ etc. Excluded was the usual choice of the direction along the axis $z$ because it turned out to be the only one which did not lead to abnormally decaying fields. The replacement of $\rho$ was done for more convenient working in Cartesian coordinates. The Riemann-Silberstein vector $\vec{F}$ was calculated according to the expression
$\vec{F}=\nabla\times\nabla\times\vec{\Pi}+\frac{i}{c}\frac{\partial}{\partial t}\nabla\times\vec{\Pi}.$ (30)
We checked that with $\vec{\Pi}\left(x,y,z,t\right)=\vec{m}f\left(x,y,z,t\right)$ the vector $\vec{F}$ indeed obeys the Maxwell equations for free space
×$\nabla\times\vec{F}-\frac{i}{c}\frac{\partial}{\partial t}\vec{F}=0,\quad\nabla\cdot\vec{F}=0.$ (31)
Real electric and magnetic fields were calculated by the known relations
$\vec{E}=\sqrt{\frac{2}{{{\varepsilon}_{0}}}}\,\Re\vec{F},\;\;\;\vec{B}=c^{-1}\sqrt{\frac{2}{{{\varepsilon}_{0}}}}\,\Im\vec{F}.$ (32)
The following results were obtained : for all possible choices of $\vec{m}$, excluding $(0,0,1)$, the $x$-components and $y$-components of electric and magnetic fields both have abnormal asymptotics when $z\rightarrow\infty$ while the $z$-components do not. Hence, $E_{x}$, $E_{y}$, $B_{x}$, and $B_{y}$ behave similarly to $f\left(\rho,z,t\right)$, see Eqs. (14), (15), (28), and (29). In other words, when propagating in the positive direction of the axis $z$ they decay as $\sim 1/(z=ct)^{3/4}$, i.e., slower than the normal decay $\sim 1/(z=ct)$. If instead of the power $3/4$ another value between $1/2$ and $1$ is used in the definition of the splash pulse, the law of the decay changes correspondingly. Fig. 7 as an example of behavior of $E_{x}$ shows that in comparison with the scalar-valued wave function in Fig. 3 the bipolar profile has become richer. This is due to the circumstance that according to Eq. (30) the electromagnetic fields are expressed from $f\left(x,y,z,t\right)$ through double temporal and spatial derivatives. It is interesting to note that only $\partial^{2}_{z,z}$ and $\partial^{2}_{t,z}$ lead to the terms with abnormal asymptotics. In Fig. 7 there is a negative peak at a location $z<0$, which is stronger than the main positive peak. There may arise a doubt as to how then the expanding pulse’s edge at the domain $z<0$ decays faster than that in the domain $z>0$. However, a version of the plot in Fig. 6 that we made for $ctE_{x}\left(0,0,z,t\right)$ demonstrates that with growing of the value of $z>0$ the peak strength increases indefinitely while its counterpart at $z<0$ does not, i.e., both behave in accordance with Eqs. (14) and (15).
In acoustics also the potential $f\left(x,y,z,t\right)$ is not directly observable. The main physical observables are the flow velocity $\vec{v}=-\nabla\,\Re f\left(x,y,z,t\right)$ and the excess pressure $p=\rho\,\partial_{t}\,\Re f\left(x,y,z,t\right)$, where $\rho$ is the mass density. Both have abnormal asymptotic behavior analogous to that of $f\left(x,y,z,t\right)$. However, like in the case of electromagnetic fields, these quantities must be consistent with the basic equations of fluid dynamics— the continuity equation and Euler’s equation of motion [27]:
$\frac{\partial\rho}{\partial t}+\nabla\cdot(\rho\,\vec{v})$ $=0\,,$ (33)
$\rho\left[\frac{\partial\vec{v}}{\partial t}+(\vec{v}\cdot\nabla)\vec{v}\,\right]$ $=-\nabla\,\rho\,.$ (34)
We checked that when calculated $\vec{v}$ and $p$ to first order, indeed they obey Eqs. (33) and (34), i.e., they are physical observables.
The fact that the fractional splash pulse is not unidirectional does not imply that its physical feasibility is impossible. Its backward-propagating components can be suppressed through appropriate adjustments of the parameters $a_{1}$ and $a_{2}$. A practical approach to addressing this issue is to generate finite-aperture approximations of the pulse, similar to what was done in the experimental realization of the focus wave mode [28, 29].
As to the primitives in Eqs. (3) and (9), the electromagnetic and acoustic physical fields derived from them possess normal asymptotics. Therefore these primitives are mainly of theoretical interest.
VI Conclusion
The solutions to the scalar wave equation with unusual asymptotic behavior in the far zone considered in this article are clearly of interest in mathematical physics. It is remarkable that there is a number of solutions which decay with propagation distance slower than the common inverse proportional law and yet have finite energy. Especially interesting is the family of fractional splash pulses. Indeed, it can be easily verified that if one changes in Eq. (13) the power $3/4$ with any other value between $1/2$ and 1, the pulse has slower than normal decay and at the same time its wave function is square-integrable and has finite total energy. These characteristics are passed on to physical electromagnetic as well as to acoustic fields derivable from the scalar wavefunctions of the fractional splash pulses. Hence, the fields with abnormally slow decay in the far zone can be in principle implemented as near-cycle pulses of radiofrequency, optical or acoustical waves.
Figure 1: The real and imaginary parts of $U\left(\rho,z,t\right)$ at the instants $ct=0$ and $ct=4$ ($c\equiv 1$). The axis $x$ represents any axis transverse to the propagation axis $z$. The sign of the imaginary part has been reversed. The pulse width parameters are chosen as follows: $ct_{s}=0.3$ and $z_{s}=0.1$.
Figure 2: ”Top view” of the 3D plot of the real part of $U\left(\rho,z,t\right)$ at four instants $t$. See caption of Fig. 1.
Figure 3: The real and imaginary parts of $f\left(\rho,z,t\right)$ at $ct=0$ and $ct=10$. At positive times, the plot has mirror symmetry $z\rightarrow-z$ The pulse width parameters are $a_{1}=1$ and $a_{2}=2$. See also caption of Fig. 1.
Figure 4: Radial dependence of the imaginary parts of the expanding and converging splash waves defined in Eq. (21) and Eq. (22) at instants $ct=30$, $ct=170$, and $ct=-100$. The first two curves have been multiplied by $-R$, the third curve by $-50R$, and the last curve by $-R$. The pulsewidth parameter $ct_{s}=2$.
Figure 5: Radial dependencies of the real parts of the waves defined in Eq. (24), Eq. (25), and Eq. (23) at instants $ct=30$ and $ct=160$. All curves have been multiplied by $R$. The pulse width parameter is $ct_{s}=2$. For comparison, a logarithmically diverging dependence is shown by the dotted curve, whereas the factor $1/2$ has been inserted in accordance with the presence of the same factor in Eqs. (24) and (25)
Figure 6: Solid curves—time dependencies of the imaginary parts (multiplied by $ct$) of the fractional splash pulse $f\left(\rho=0,z,t\right)$ at points $z=30$, $z=90$, and $z=150$; dashed curve—an analogue of the expression under $lim$ in Eq. (2), i.e., $ct\cdot\Im f\left(\rho=0,z=ct+\Delta,t\right)$. For comparison, the dotted curve shows a diverging dependence $\sin(3\pi/8)\sqrt[4]{2ct}/2a_{1}$. The pulse width parameters are $a_{1}=3$ and $a_{2}=2$. $\Delta=-3/4$. The two last curves are shown only for the region $ct>0$.
Figure 7: The $x$-component of the real electric field $E_{x}\left(x,y=0,z,t\right)$ at $ct=10$ calculated from the Hertz vector (1,1,0)$f\left(\rho,z,t\right)$. The pulse width parameters are $a_{1}=1$ and $a_{2}=2$. Cf., also, with Fig. 3.

### v2

@@
 The former two have normal asymptotics in the sense of finite values of the limit of the type in Eq.
-(2) as they have to be since both consist of derivatives of $U\left(\rho,z,t\right)$.
+(2) as they have to be since both consist of derivatives of $U\left(\rho,z,t\right)$ which have normal asymptotics.
 Despite the unidirectionality of $U\left(\rho,z,t\right)$—integration as a linear operation preserves the unidirectional spectrum of $u\left(\rho,z,t\right)$— in certain spatial regions the energy flux turned out to be directed opposite to the azis $z$, and at some points the velocity of such an energy backflow even reached almost the value $-c$.
@@
 Even if one modifies Eq.
-(17) by replacements $\rho=ct/\sqrt{2}+\Delta,\,z=ct/\sqrt{2}+\Delta$ in order to keep the point (running along the diagonal direction) close to the pulse’s top, the expression for the limit remains finite.
+(17) by the replacements $\rho=ct\,\sin\alpha+\Delta,\,z=ct\,\cos\alpha+\Delta$, where $\alpha\approx 0$, in order to direct the limit taking almost along the axis $z$, the expression of the limit remains finite.
+This is explained by the relative narrowing of the transverse width of the pulse with the propagation distance $z=ct$ and the remarkable feature that only the top of the pulse decays abnormally with distance.
+Specifically, if we calculate in 3D the solid angle $\Omega=\pi(HWHM)^{2}/(ct)^{2}$ formed by a circular area with radius equal to transverse half width at half maximum (HWHM) of the pulse peak (of the real part) in the expanding sphere of radius $ct$, then we get the following numerical results.
+When $ct=10$ as in Fig.
+3, then $\Omega\approx 2\,sr$; when $ct=100$, then $\Omega\approx 0.2\,$; when $ct=1000$, then $\Omega\approx 0.02\,sr$, and when $ct=10000$, then $\Omega\approx 0.002\,sr$.
+Hence, although the value of the HWHM increases with propagation, the angular half-width of the peak of the pulse decreases 10-fold if the propagation distance increases 10-fold and at infinity turns to zero around the axis $z$.
 We calculated the Poynting vector and energy density of $f\left(\rho,z,t\right)$.
@@
 Fig.
-7 as an example of behavior of $E_{x}$ shows that in comparison with the scalar-valued wave function in Fig.
-3 the bipolar profile has become richer.
+7 illustrating the behavior of $E_{x}$ shows that in comparison with the scalar-valued wave function in Fig.
+3, the pulse’s maximum has become more salient, and the solid angle formed by a circle with radius equal to HWHM of the pulse is more than six times smaller than that of the scalar pulse resulting in $\Omega\approx 0.0003\,sr$ if $ct=10000$ (see Section III).
 This is due to the circumstance that according to Eq.
-(30) the electromagnetic fields are expressed from $f\left(x,y,z,t\right)$ through double temporal and spatial derivatives.
-It is interesting to note that only $\partial^{2}_{z,z}$ and $\partial^{2}_{t,z}$ lead to the terms with abnormal asymptotics.
-In Fig.
-7 there is a negative peak at a location $z<0$, which is stronger than the main positive peak.
-There may arise a doubt as to how then the expanding pulse’s edge at the domain $z<0$ decays faster than that in the domain $z>0$.
-However, a version of the plot in Fig.
-6 that we made for $ctE_{x}\left(0,0,z,t\right)$ demonstrates that with growing of the value of $z>0$ the peak strength increases indefinitely while its counterpart at $z<0$ does not, i.e., both behave in accordance with Eqs.
-(14) and (15).
+(30) the electromagnetic fields are expressed from $f\left(x,y,z,t\right)$ through double temporal and spatial derivatives which take their highest values in the region of the pulse maximum.
+It is interesting to note that only $\partial^{2}_{z,z}$ and $\partial^{2}_{t,z}$ lead to terms with abnormal asymptotics.
 In acoustics also the potential $f\left(x,y,z,t\right)$ is not directly observable.
@@
 Its backward-propagating components can be suppressed through appropriate adjustments of the parameters $a_{1}$ and $a_{2}$.
-A practical approach to addressing this issue is to generate finite-aperture approximations of the pulse, similar to what was done in the experimental realization of the focus wave mode [28, 29].
-As to the primitives in Eqs. (3) and (9), the electromagnetic and acoustic physical fields derived from them possess normal asymptotics. Therefore these primitives are mainly of theoretical interest.
+A practical approach to address this issue is to generate unidirectional finite-aperture approximations of the pulse, similar to what was done in the experimental realization of the focus wave mode [28, 29].
+As to the primitives in Eqs. (3) and (9), the electromagnetic and acoustic physical fields derived from them possess normal asymptotics. Therefore, these primitives are mainly of theoretical interest.
 VI Conclusion
 The solutions to the scalar wave equation with unusual asymptotic behavior in the far zone considered in this article are clearly of interest in mathematical physics.
-It is remarkable that there is a number of solutions which decay with propagation distance slower than the common inverse proportional law and yet have finite energy.
+It is remarkable that there are several solutions that decay with propagation distance more slowly than the common inverse proportional law, yet still possess finite energy.
 Especially interesting is the family of fractional splash pulses.
 Indeed, it can be easily verified that if one changes in Eq.
-(13) the power $3/4$ with any other value between $1/2$ and 1, the pulse has slower than normal decay and at the same time its wave function is square-integrable and has finite total energy.
+(13) the power $3/4$ with any other value between $1/2$ and 1, the pulse has a slower than normal decay and, at the same time, its wave function is square-integrable and has finite total energy.
 These characteristics are passed on to physical electromagnetic as well as to acoustic fields derivable from the scalar wavefunctions of the fractional splash pulses.
-Hence, the fields with abnormally slow decay in the far zone can be in principle implemented as near-cycle pulses of radiofrequency, optical or acoustical waves.
+Hence, the fields with abnormally slow decay in the far zone can in principle be implemented as near-cycle pulses of radiofrequency, optical, or acoustical waves.
 Figure 1: The real and imaginary parts of $U\left(\rho,z,t\right)$ at the instants $ct=0$ and $ct=4$ ($c\equiv 1$).

### v3

@@
 T.
-Wu introduced models of “electromagnetic missiles” whose decay could be made arbitrarily slower than the usual inverse distance by an appropriate choice of the high frequency portion of the source spectrum.
-Very recent work by Plachenov and Kiselev introduced a finite-energy scalar wave solution, different from Wu’s, decaying slower than inversely proportional with the distance.
+Wu [J.
+Appl.
+Phys.
+57, 2370 (1985)] introduced models of "electromagnetic missiles" whose decay could be made arbitrarily slower than the usual inverse distance by an appropriate choice of the high frequency portion of the source spectrum.
+Very recent work by Plachenov and Kiselev [Diff.
+Eqs.
+60, 1634 (2024).] introduced a finite-energy scalar wave solution, different from Wu’s, decaying slower than inversely proportional with the distance.
 A physical explanation for the unusual asymptotic behavior of the latter will be given in this article.
@@
 A proof of feasibility of acoustic and electromagnetic fields with the abnormal asymptotics will be described.
-wave equation, unidirectional wavepackets, ultrashort electromagnetic and acoustical pulses
-preprint: APS/123-QED
-I Introduction
+journal: opticajournalarticletype: Research Article
+1 Introduction
 It is a textbook truth that solutions of the wave equation decay inversely proportional to distance in the far field or the wave zone (see, e.g., [1, 2]).
-However, a long time ago Wu [3] showed that the decay can be made arbitrarily slower by a careful choice of the high-frequency wing of the source spectrum.
-In the 1980s, closely related to Wu’s ”electromagnetic missiles”, another subject dealing with uncommon solutions of the wave equation emerged—the so-called localized or non-diffracting or space-time wave packets possessing exotic properties, such as self-healing, superluminality, and invariant propagation without spread or decay, theoretically up to infinity.
+Long time ago Wu [3] showed that the decay of radiatation from a finite-size antenna can be made arbitrarily slower thanks to stretching the Fresnel region for a part of the source spectrum by a careful choice of the high-frequency wing of the spectrum.
+However, subsequent experiments did not substantiate claims of unlimited slow decay [4].
+A particular study of quasi-missile behavior was undertaken by Shaarawi et al.
+[5].
+They examined in detail the asymptotic behavior of a particular scalar localized wave, referred to as double-exponential pulse (DEX).
+In the 1980s, closely related to Wu’s "electromagnetic missiles", another subject dealing with uncommon solutions of the wave equation emerged—the so-called localized or non-diffracting or space-time wave packets possessing exotic properties, such as self-healing, superluminality, and invariant propagation without spread or decay, theoretically up to infinity.
 By now this subject has grown into a large research field of its own, see monograph [6] and reviews [7, 8, 9, 10, 11].
 Therefore, the decay in the wave zone as a specific subject in mathematical physics has remained somewhat in the shadows.
-However, in recent years, several papers on the study of asymptotical behavior have appeared, see [10, 11, 12] and Refs.
-therein.
-The very recent paper by Plachenov and Kiselev [12] introduced a finite-energy solution, different from Wu’s model [3], that decays slower than inversely proportional to the distance.
-They consider the following simple solution of the homogeneous wave equation (we have somewhat changed the designations)
+However, in recent years, several papers on the study of asymptotical behavior have appeared, see [12, 13, 14] and references therein.
+The very recent paper by Plachenov and Kiselev [14] (see also [15]) introduced a finite-energy solution, different from Wu’s model [3], that decays slower than inversely proportional to the distance.
+They consider the following simple solution
 $\psi\left(R,t\right)=\frac{1}{\left(ct+ict_{s}\right)^{2}-R^{2}}\;,$ (1)
-where $R=\left(x^{2}+y^{2}+z^{2}\right)^{1/2}$ is the distance of the field point from the origin, $c$ is the velocity of light (or sound), and $t_{s}$ is a constant that determines the width of the pulse.
+to the homogeneous wave equation
+$\frac{1}{\rho}\frac{\partial}{\partial\rho}\left(\rho\frac{\partial\psi}{\partial\rho}\right)+\frac{\partial^{2}}{\partial z^{2}}\psi=\frac{\partial^{2}}{\partial(ct)^{2}}\psi\;,$ (2)
+where $R=\left(\rho^{2}+z^{2}\right)^{1/2}$ is the distance of the field point from the origin, $c$ is the velocity of light (or sound), and $t_{s}$ is a constant that determines the width of the pulse.
+Here differently from [14] we have used somewhat changed designations and wrote the wave equation in cylindrical coordinates for the axisymmetric case which is satisfied by the solutions introduced in the following sections.
 $\psi\left(R,t\right)$ is a special case of a splash pulse [16].
-It is frequently used for modeling ultrashort pulses and constitutes a spherically propagating pulse, while the spherical surface, where the pulse peaks at $R\approx ct,$ collapses into the origin at negative times $t<0$ and expands for $t>0.$ The real part $\operatorname{Re}\psi\left(R,t\right)$ represents a bipolar pulse and the imaginary part $\operatorname{Im}\psi\left(R,t\right)$ a unipolar one everywhere where $R\gg ct_{s}$, and vice versa near the origin where $R<ct_{s}$.
+It is frequently used to model ultrashort pulses and constitutes a spherically propagating pulse, while the spherical surface, where the pulse peaks at $R\approx ct,$ collapses into the origin at negative times $t<0$ and expands for $t>0.$ The real part $\operatorname{Re}\psi\left(R,t\right)$ represents a bipolar pulse and the imaginary part $\operatorname{Im}\psi\left(R,t\right)$ a unipolar one when the pulse moves in the far field where $R\gg ct_{s}$, and vice versa near the origin where $R<ct_{s}$.
 They consider the limit
@@
 The limit of the type of Eq.
-(2) plays an important role in various fields of mathematical physics and it is known that for finite-energy ”well-behaving” solutions of the wave equation the limit $L$ should be finite (incl.
+(3) plays an important role in various fields of mathematical physics and it is known that for finite-energy "well-behaving" solutions of the wave equation the limit $L$ should be finite (incl.
 $L=0$), see, e.g., [1, 2, 12, 13].
 The authors of [14] showed that for the primitive (antiderivative) of $\psi\left(R,t\right)$
-[…]
-in contradistinction to the usual behaviour of fields in the wave zone, the quantity in Eq. (2) is diverging:
+$\begin{split}\Psi\left(R,t\right)=\int_{-\infty}^{t}\psi\left(R,t^{\prime}\right)dt^{\prime}&=\\ =\frac{1}{2cR}\ln\frac{ct+ict_{s}-R}{ct+ict_{s}+R},\end{split}$ (4)
+which of course also obeys the wave equation, but in contradistinction to the usual behaviour of fields in the wave zone, the quantity in Eq. (3) is diverging:
 $L=-\left(1/2c\right)\lim_{t\rightarrow\infty}\ln ct=-\infty.$ (5)
-They also showed that $\Psi\left(R,t\right)$ has finite total energy, i.e., it is not an unphysical solution.
-The purpose of our article is to present additional examples of pulses with unusual asymptotics and to reveal physical causes of abnormally slow decay.
-In the next section we consider one of such examples—the primitive (antiderivative) of the wave function of a unidirectional pulse.
-Sect.
-III is devoted to another example—the fractional splash pulse which already itself, without taking its time integral, decays very slowly in the far zone.
+They also showed that $\Psi\left(R,t\right)$ has finite total energy, i.e., may not be an unphysical solution.
+However, both the scalar field energy density and Poynting vector have normal asymptotics in the sense of finite values of the limit of the type in Eq.
+(3).
+The general motivation of our article is to draw attention to the phenomenon of abnormally slow decay, which is quite unexpected, at least in the case of physically feasible waves.
+Our aim is to find additional examples of wave pulses with unusual asymptotics, and to reveal the causes of their abnormally slow decay.
+We pay special attention to determining the finiteness of the energy of the found pulses because this property is a crucial one for their physical feasibility In the next section we consider one of such examples—the primitive (antiderivative) of the wave function of a unidirectional pulse.
+Section 3 is devoted to another example—the fractional splash pulse which already itself, without taking its time integral, decays very slowly in the far zone.
 The authors in the short communication [14] neither illustrate graphically the behavior of $\Psi\left(R,t\right)$ nor interpret physically its unusual asymptotics.
-In Sec.
-IV we try to fill in this gap.
-In Sect.
-V we discuss some subtleties of the slow decay of the other pulses studied in Sect.
-II and III and consider the physical feasibility in optics, electromagnetics, or acoustics of the pulses considered in this study.
-II Hemispherical unidirectional pulse
+In Section 4 we try to fill in this gap.
+In Section 5 we discuss some subtleties of the slow decay of the other pulses studied in Sections 2 and 3 and consider the physical feasibility in optics, electromagnetics, or acoustics of the pulses considered in this study.
+2 Hemispherical unidirectional pulse
 Now we will consider a cylindrically symmetric pulse which is unidirectional in the sense that all its plane-wave constituents propagate solely into the hemisphere with the axial component of the wave vector $k_{z}>0$.
@@
 Similarly to $\psi\left(R,t\right)$ in Eq.
-(1), the real and imaginary parts of $u\left(R,t\right)$ represent bipolar and unipolar pulses, but in contrast to the spherically symmetric case, in the given case the pulses ”ride” on a collapsing-expanding tube-like surface with a hemispherical surface inside (see plots below).
+(1), the real and imaginary parts of $u\left(R,t\right)$ represent bipolar and unipolar pulses, but in contrast to the spherically symmetric case, in the given case the pulses "ride" on a collapsing-expanding tube-like surface with a hemispherical surface inside.
 $u\left(\rho,z,t\right)$ has normal asymptotics:
@@
 But how does its primitive behave?
-Taking the integral of $u\left(\rho,z,t\right)$ is a complicated task.
-It can be found neither in tables of integrals nor can be taken by the newest versions of computation packages Mathematica and Maple.
-The integral over time from $\infty$- ∞ to $t$ was taken first by setting $t_{s}=z_{s}=0$ and then making replacements $z\rightarrow z-iz_{s}$ and $t\rightarrow t+it_{s}$.
+The integral over time from -$\infty$ to $t$ was taken first by setting $t_{s}=z_{s}=0$ and then making replacements $z\rightarrow z-iz_{s}$ and $t\rightarrow t+it_{s}$.
 After some algebra, we found the following expression for the primitive, where $g\left(\rho,t\right)$ is given in Eq.
@@
 $U\left(\rho,z,t\right)$ $=\frac{1}{2h\left(\rho,z\right)}\left[U_{1}\left(\rho,z,t\right)+U_{2}\left(\rho,z,t\right)\right],$ (10)
-$U_{1}\left(\rho,z,t\right)$ $\equiv\ln\frac{ct+ict_{s}+h\left(\rho,z\right)}{ct+ict_{s}-h\left(\rho,z\right)},\qquad h\left(\rho,z\right)\equiv\sqrt{\rho^{2}+\left(z-iz_{s}\right)^{2}},$
-$U_{2}\left(\rho,z,t\right)$ $\equiv\ln\frac{\left[\left(-ict+ct_{s}\right)\left(z-iz_{s}\right)+g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z-iz_{s}-h\left(\rho,z\right)\right]}{\left[\left(-ict+ct_{s}\right)\left(z-iz_{s}\right)-g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z-iz_{s}+h\left(\rho,z\right)\right]}.$
-The behavior of $U\left(\rho,z,t\right)$ is depicted in Figs.
-1 and 2.
+$U_{1}\left(\rho,z,t\right)$ $\equiv\ln\frac{ct+ict_{s}+h\left(\rho,z\right)}{ct+ict_{s}-h\left(\rho,z\right)},$
+$h\left(\rho,z\right)\equiv\sqrt{\rho^{2}+z^{*2}},$
+$U_{2}\left(\rho,z,t\right)$ $\equiv\ln\frac{\left[ct^{*}z^{*}+g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z^{*}-h\left(\rho,z\right)\right]}{\left[ct^{*}z^{*}-g\left(\rho,t\right)h\left(\rho,z\right)\right]\left[z^{*}+h\left(\rho,z\right)\right]},$
+where $ct^{*}$ stands for $ict+ct_{s}$ and $z^{*}$ for $z-iz_{s}$. The behavior of $U\left(\rho,z,t\right)$ is depicted in Figs. 1 and 2.
+Figure 1: The real and imaginary parts of $U\left(\rho,z,t\right)$ at the instants $ct=0$ and $ct=4$ ($c\equiv 1$).
+As the plot is axisymmetric, the axis $x$ represents any axis transverse to the propagation axis $z$ and, in distinction from the radial coordinate $\rho,$ also takes negative values.
+The sign of the imaginary part has been reversed.
+The pulse width parameters are chosen as follows: $ct_{s}=0.3$ and $z_{s}=0.1$.
+Figure 2: "Top view" of the 3D plot of the real part of $U\left(\rho,z,t\right)$ at four instants $t$. See caption of Fig. 1.
 We see that the pulse peaks on the surface of a collapsing-expanding tube with a hemisphere inside it, while the radius of both is equal to $|ct|$.
@@
 $\lim_{t\rightarrow\infty}\left[ct\,U\left(\rho,z,t\right)\right]_{\rho=ct+\Delta}=\infty~,$ (13)
-Another exception where the limit proves to be finite and equals to $\ln 2$ is when time $t$ runs to $\infty$- ∞ but $z$ to +$\infty$+ ∞.
+Another exception where the limit proves to be finite and equals to $\ln 2$ is when time $t$ runs to -$\infty$ but $z$ to $\infty$.
 In both these exceptional cases the pulse peak is not involved into the process, as can be seen by looking at Figs.
 1 and 2.
-We calculated the Poynting vector and energy density of $U\left(\rho,z,t\right)$ according to known formulas (see, e.g., Eqs.
+We calculated the Poynting vector and energy density of $U\left(\rho,z,t\right)$ according to known formulas [23] (see also Eqs.
 (1)-(5) of Ref.
@@
 The latter is finite.
-The former two have normal asymptotics in the sense of finite values of the limit of the type in Eq.
+However, both the Poynting vector and the energy density have normal asymptotics in the sense of finite values of the limit of the type in Eq.
 (3) as they have to be since both consist of derivatives of $U\left(\rho,z,t\right)$ which have normal asymptotics.
-Despite the unidirectionality of $U\left(\rho,z,t\right)$—integration as a linear operation preserves the unidirectional spectrum of $u\left(\rho,z,t\right)$— in certain spatial regions the energy flux turned out to be directed opposite to the azis $z$, and at some points the velocity of such an energy backflow even reached almost the value $-c$.
+Despite the unidirectionality of $U\left(\rho,z,t\right)$—integration as a linear operation preserves the unidirectional spectrum of $u\left(\rho,z,t\right)$— in certain spatial regions the energy flux turned out to be directed opposite to the azis $z$, and at some points the velocity of such an energy backflow even reached almost the value c-c.
 We calculated the electric and magnetic fields from $U\left(\rho,z,t\right)$ using the Hertz vector technique and proved that they have normal asymptotics.
 This is understandable, as the technique involves taking derivatives of the scalar field.
-Finally, we found that the imaginary part of $U\left(\rho,z,t\right)$ is a ”strange” field [20, 21, 22] in the sense that in certain regions the time integral from $\infty$- ∞ to +$\infty$+ ∞ is not zero.
+Finally, we found that the imaginary part of $U\left(\rho,z,t\right)$ is a "strange" field [24, 25, 26] in the sense that in certain regions the time integral from -$\infty$ to $\infty$ is not zero.
 This is understandable, because the imaginary part is a unipolar pulse.
-III Fractional splash pulse
-Members of a general family of splash pulses are derived as spectral superpositions of the most investigated localized propagation-invariant unipolar pulse—the focus wave mode [5, 6, 4].
+3 Fractional splash pulse
+It is known that specific members of a general family of splash pulses are derivable as spectral superpositions of the most investigated localized propagation-invariant unipolar pulse—the focus wave mode [7, 8, 6, 16].
 A specific fractional splash pulse is given by the following expression
 $f\left(\rho,z,t\right)$ $=\frac{1}{a_{1}+i\left(z-ct\right)}$ (14)
-×$\times\frac{1}{\left(a_{2}-i\left(z+ct\right)+\rho^{2}\left(a_{1}+i\left(z-ct\right)\right)^{-1}\right)^{3/4}},$
+$\times\frac{1}{\left(a_{2}-i\left(z+ct\right)+\rho^{2}\left(a_{1}+i\left(z-ct\right)\right)^{-1}\right)^{3/4}},$
 where $a_{1}$ and $a_{2}$ are positive parameters.
@@
 The fractional pulse above corresponds to $\nu=-1/4$.
+We note that in the case of DEX pulses [5], which are presented as differences of two splash pulses, the slow decay appears also if $\nu<-1$, but such pulses are not square integrable and the total energy of corresponding fields diverges.
 The behavior of $f\left(\rho,z,t\right)$ is depicted in Fig.
@@
 The real part has mirror symmetry with respect to time reversal $t\rightarrow-t$, while the imaginary part additionally reverses its sign.
-Outside the region of the origin, the pulse tops of the real part and the imaginary part do not coincide: the top of the former is shifted farther from the origin $z=0$ than the point $z=ct$ by a generally small distance about $32a_{1}/$ 3, and the top of the imaginary part is shifted closer by the distance about $a_{1}/4$.
+Outside the region of the origin, the pulse tops of the real part and the imaginary part do not coincide: the top of the former is shifted farther from the origin $z=0$ than the point $z=ct$ by a generally small distance about $2a_{1}/3$, and the top of the imaginary part is shifted closer by the distance about $a_{1}/4$.
 This follows from a detailed inspection of Eq.
-(13), as well as from plots which will be presented in Section V.
+(14), as well as from plots which will be presented in Section 5.
 But the most remarkable distinction from the pulses considered in the previous sections is that the fractional splash pulse itself, without the need to take its antiderivative, exhibits abnormal asymptotics as the limits of the type in Eq.
@@
 (15) shows that the peaks of both the real and imaginary parts decay slower than the normal decay $\sim 1/z$ when one moves with the pulse top along the positive direction of the axis $z$ towards infinity.
-How they approach infinity will be discussed in Sect.
-V.
+How they approach infinity will be discussed in Section 5.
 As was mentioned above, the tops of the pulses of the real and imaginary parts of $f\left(\rho,z,t\right)$ are shifted from the point $z=ct$, and therefore the value of $\Delta$ affects differently the limits of the parts.
@@
 (15) shows that at this point both $ct\,\Re f\left(\rho,z=ct,t\right)$ and $ct\,\Im f\left(\rho,z=ct,t\right)$ approach infinity in the same way.
+Figure 3: The real and imaginary parts of $f\left(\rho,z,t\right)$ at $ct=0$ and $ct=10$.
+At positive times, the plot has mirror symmetry $z\rightarrow-z$ The pulse width parameters are $a_{1}=1$ and $a_{2}=2$.
+See also caption of Fig.
+1.
 Eqs.
@@
 The total energy is finite.
-$f\left(\rho,z,t\right)$ is square integrable, which is also important from a physical point of view as will be discussed in Sect.
-V.
+$f\left(\rho,z,t\right)$ is square integrable (see Appendix), which is also important from a physical point of view as will be discussed in Section 5.
 We studied also other fractional splash pulses which are given by the same expression in Eq.
@@
 If the power is $1/2$, the pulse has the same properties as $f\left(\rho,z,t\right)$, but unfortunately its total energy is infinite.
-But if the value of the power is between $1/2$ and $1$, i.e., if $\frac{1}{2}<\nu+1<1$, the pulse has slower than normal decay and at the same time its wave function is square-integrable and has finite total energy.
-IV Causes of uncommon asymptotical behavior at infinity
+But if the value of the power is between $1/2$ and $1$, i.e., if $\frac{1}{2}<\nu+1<1$, the pulse has slower than normal decay and at the same time its wave function is square-integrable and has finite total energy (see Appendix).
+4 Causes of uncommon asymptotical behavior at infinity
 It is known that solutions to the homogeneous wave equation can be expressed by a convolution of the density of a fictitious Huygens source (coupled with sink) and the free-space propagator or the Riemann - Schwinger function
@@
 where $G_{+}$ and $G_{-}$ are the retarded and advanced Green functions, respectively. By doing this with the delta-point-like source function
-$\rho(\textbf{r},t)=\delta(\textbf{r})\hskip 2.0pt\frac{1}{2c(t+it_{s})}~{},$ (19)
+$\rho(\textbf{r},t)=\delta(\textbf{r})\hskip 2.0pt\frac{1}{2c(t+it_{s})}~,$
 a field like $\psi\left(R,t\right)$ in Eq.
@@
 Here $\psi_{-}$ represents the spherical pulse converging at negative times to the origin and $\psi_{+}$ expanding from it at positive times.
+Figure 4: Radial dependence of the imaginary parts of the expanding and converging waves defined in Eq.
+(21) and Eq.
+(22) at instants $ct=30$, $ct=170$, and ct=100ct=-100.
+All curves have been multiplied by R-R and the third curve additionally by 50.
+The pulsewidth parameter $ct_{s}=2$.
 In Fig.
@@
 (22), respectively, are plotted against the radius $R$ at three different instants of time.
-The curves are plotted with reverse sign for better readability and are scaled—multiplied by $R$—in order to eliminate the factor $1/R$.
+The curves are plotted with reverse sign for better readability and are scaled—multiplied by RR—in order to eliminate the factor $1/R$.
 Thanks to the scaling we observe that the peaks of the pulses at different time instants are of equal height.
 This correponds to finitness of the limit in Eq.
-(2), and means that the splash pulse has indeed normal asymptotics, as mentioned already in Sect.1.
+(3), and means that the splash pulse has indeed normal asymptotics, as mentioned already in Section 1.
 The same holds also for the real parts of the wave functions in Eqs.
@@
 (20)-(22), the primitive of the splash pulse $\Psi\left(R,t\right)$ can be decomposed into expanding and converging waves, as can be seen readily from Eq.
-(3)  [^1] If one integrates $\psi_{+}\left(R,t\right)$ and $\psi_{-}\left(R,t\right)$ over time from $\infty$- ∞ up to any finite value $t$, then the real parts of both $\Psi_{+}\left(R,t\right)$ and $\Psi_{-}\left(R,t\right)$ acquire an additive integration constant $C=-ln(\infty)$ which cancels out in Eq.
+(4)  [^1] If one integrates $\psi_{+}\left(R,t\right)$ and $\psi_{-}\left(R,t\right)$ over time from -$\infty$ up to any finite value $t$, then the real parts of both $\Psi_{+}\left(R,t\right)$ and $\Psi_{-}\left(R,t\right)$ acquire an additive integration constant $C=-ln(\infty)$ which cancels out in Eq.
 (23) and therefore does not affect $\Psi\left(R,t\right)$., viz.:
@@
 (23), are plotted against the radius $R$ at two different instants of time.
-We no longer deal with the imaginary parts because their asymptotics is not abnormal.
-This is understandable, because as integrands in Eq.
+We no longer deal with the imaginary parts because their asymptotics are not abnormal.
+This is understandable because, as integrands in Eq.
 (4), the imaginary parts of $\psi_{+}\left(R,t\right)$ and $\psi_{-}\left(R,t\right)$ decay with $t\rightarrow\infty$ as $t^{-2}$, while the real parts as $t^{-1}$.
-The curves are multiplied by $R$ and so become scaled appropriately for their comparison.
+Figure 5: Radial dependencies of the real parts of the waves defined in Eq.s (24), (25), and (23) at instants $ct=30$ and $ct=160$.
+All curves have been multiplied by $R$.
+The pulse width parameter is $ct_{s}=2$.
+For comparison, a logarithmically diverging dependence is shown by the dotted curve, whereas the factor $1/2$ has been inserted in accordance with the presence of the same factor in Eqs.
+(24) and (25).
+The curves are multiplied by $R$ and thus are appropriately scaled for comparison.
 We observe that the peaks of the real part of $\Psi_{+}(R,t)$ at different time instants are of equal height.
-This corresponds to finitness of the limit in Eq.(2) and means that the expanding part of the primitive, taken separately, has normal asymptotics.
-But we also observe, that if the tails of the converging waves are added— according to Eq.
-(23) with negative sign—the peaks of $R\cdot\Re\left[\Psi\left(R,t\right)\right]$ of the splash pulse grow logarithmically toward $\infty$- ∞.
-Indeed, the sum of the ordinate of the peak at $R=30$ of $R\cdot\Re\left[\Psi_{+}(R,30)\right]$ and that of $-R\cdot\Re\left[\Psi_{-}(R,30)\right]$ is exactly equal to the ordinate of the peak of $R\cdot\Re\left[\Psi(R,30)\right]$ at the same abscissa $R=30$.
+This corresponds to the finitness of the limit in Eq.(3) and means that the expanding part of the primitive, taken separately, has normal asymptotics.
+But we also observe that if the tails of the converging waves are added— according to Eq.
+(23) with negative sign—the peaks of $R\cdot\Re\left[\Psi\left(R,t\right)\right]$ of the splash pulse grow logarithmically toward -$\infty$.
+Indeed, the sum of the ordinate of the peak at $R=30$ of $R\cdot\Re\left[\Psi_{+}(R,30)\right]$ and that of $R\cdot\Re\left[\Psi_{-}(R,30)\right]$ is exactly equal to the ordinate of the peak of $R\cdot\Re\left[\Psi(R,30)\right]$ at the same abscissa $R=30$.
 The same can be observed for the curves for the time instant $ct=160$.
@@
 (3) is finite, as it has to be since the Poynting vector consists of derivatives of $\Psi\left(R,t\right)$.
-V Discussion and study of physical feasibility
+5 Discussion and study of physical feasibility
 The real part of the primitive of the unidirectional pulse has according to Eq.
@@
 Hence, all the results for the primitive of the splash pulse also apply to the radial evolution of the primitive of the unidirectional pulse.
+Figure 6: Solid curves—time dependencies of the imaginary parts (multiplied by $ct$) of the fractional splash pulse $f\left(\rho=0,z,t\right)$ at points $z=30$, $z=90$, and $z=150$; dashed curve—an analogue of the expression under limlim in Eq.
+(3), i.e., $ct\cdot\Im f\left(\rho=0,z=ct+\Delta,t\right)$.
+For comparison, the dotted curve shows a diverging dependence $\sin(3\pi/8)\sqrt[4]{2ct}/2a_{1}$.
+The pulse width parameters are $a_{1}=3$ and $a_{2}=2$.
+$\Delta=-3/4$.
+The two last curves are shown only for the region $ct>0$.
 The time dependence of the imaginary part of the fractional splash pulse $f\left(\rho,z,t\right)$ at different points on the axis $z$ is depicted in Fig.
@@
 The real part behaves analogously.
-Contrary to the case of $z-$dependence with fixed time instants, in the case of $ct-$dependence with fixed locations $z$ the tops of the curves of the imaginary parts are shifted to the right (if $t>0$) from the point where $ct$ equals to a fixed value of $z$.
+Contrary to the case of $z$-dependence with fixed time instants, in the case of ctct-dependence with fixed locations $z$ the tops of the curves of the imaginary parts are shifted to the right (if $t>0$) from the point where $ct$ equals to a fixed value of $z$.
 That is why in order for the dashed curve to touch the tops of the peaks, in the expression of $ct\cdot\Im f\left(\rho=0,z=ct+\Delta,t\right)$ the value of $\Delta$ was set to $\Delta=-a_{1}/4$.
@@
 Therefore, we could not make use of the superpositional relation between the focus wave mode and the fractional splash pulse for decomposing the latter.
-Not only the total energy of the scalar field $f\left(\rho,z,t\right)$ is finite, but also the total energy of an electromagnetic field which can be proved by a relation [26] between the square integrability of the spectrum of a splash pulse and the total energy.
+Not only the total energy of the scalar field $f\left(\rho,z,t\right)$ is finite, but also the total energy of an electromagnetic field which we proved by numerical double integration and analytical calculations with the help of spectral representation of fractional splash pulses (see Appendix).
 We undertook a thorough proof of the physical feasibility of electromagnetic fields derivable from the fractional splash pulse $f\left(\rho,z,t\right)$ with the help of a Hertz and a Riemann-Silberstein vector technique.
@@
 We checked that with $\vec{\Pi}\left(x,y,z,t\right)=\vec{m}f\left(x,y,z,t\right)$ the vector $\vec{F}$ indeed obeys the Maxwell equations for free space
-×$\nabla\times\vec{F}-\frac{i}{c}\frac{\partial}{\partial t}\vec{F}=0,\quad\nabla\cdot\vec{F}=0.$ (31)
+$\nabla\times\vec{F}-\frac{i}{c}\frac{\partial}{\partial t}\vec{F}=0,\quad\nabla\cdot\vec{F}=0.$ (31)
 Real electric and magnetic fields were calculated by the known relations
@@
 The following results were obtained: for all possible choices of $\vec{m}$, excluding $(0,0,1)$, the $x$-components and $y$-components of electric and magnetic fields both have abnormal asymptotics when $z\rightarrow\infty$ while the $z$-components do not.
-Hence, $E_{x}$, $E_{y}$, $B_{x}$, and $B_{y}$ behave similarly to $f\left(\rho,z,t\right)$, see Eqs.
-(14), (15), (28), and (29).
+Hence, $E_{x}$, $E_{y}$, $B_{x}$, and $B_{y}$ behave similarly to $f\left(\rho,z,t\right)$, see Eq.s (15), (16), (28), and (29).
 In other words, when propagating in the positive direction of the axis $z$ they decay as $\sim 1/(z=ct)^{3/4}$, i.e., slower than the normal decay $\sim 1/(z=ct)$.
 If instead of the power $3/4$ another value between $1/2$ and $1$ is used in the definition of the splash pulse, the law of the decay changes correspondingly.
+[…]
 Fig.
 7 illustrating the behavior of $E_{x}$ shows that in comparison with the scalar-valued wave function in Fig.
-3, the pulse’s maximum has become more salient, and the solid angle formed by a circle with radius equal to HWHM of the pulse is more than six times smaller than that of the scalar pulse resulting in $\Omega\approx 0.0003\,sr$ if $ct=10000$ (see Section III).
+3, the pulse’s maximum has become more salient, and the solid angle formed by a circle with radius equal to HWHM of the pulse is more than six times smaller than that of the scalar pulse resulting in $\Omega\approx 0.0003\,sr$ if $ct=10000$ (see Section 3).
 This is due to the circumstance that according to Eq.
@@
 $\rho\left[\frac{\partial\vec{v}}{\partial t}+(\vec{v}\cdot\nabla)\vec{v}\,\right]$ $=-\nabla\,\rho\,.$ (34)
-We checked that when calculated $\vec{v}$ and $p$ to first order, indeed they obey Eqs. (33) and (34), i.e., they are physical observables.
+We checked that when $\vec{v}$ and $p$ are calculated to first order, indeed they obey Eq.s (33) and (34), i.e., they are physical observables.
 The fact that the fractional splash pulse is not unidirectional does not imply that its physical feasibility is impossible.
 Its backward-propagating components can be suppressed through appropriate adjustments of the parameters $a_{1}$ and $a_{2}$.
-A practical approach to address this issue is to generate unidirectional finite-aperture approximations of the pulse, similar to what was done in the experimental realization of the focus wave mode [28, 29].
-As to the primitives in Eqs. (3) and (9), the electromagnetic and acoustic physical fields derived from them possess normal asymptotics. Therefore, these primitives are mainly of theoretical interest.
-VI Conclusion
-The solutions to the scalar wave equation with unusual asymptotic behavior in the far zone considered in this article are clearly of interest in mathematical physics.
+A practical approach to generate a unidirectional finite-aperture approximation of the pulse is using the Huygens approach [5], or following a method similar to what was done in the experimental realization of the focus wave mode [30, 31].
+As pointed out in [5], the resulting finite-energy unidirectional fractional splash mode pulse may not retain the abnormal decay along the z-direction as $z\rightarrow\infty$.
+There will be, however, an intermediate zone, roughly up to the Rayleigh distance , where the pulse will be characterized by slow decay before it assumes the usual $1/z$ behavior in the far field.
+If the aperture radius $r_{a}$ is equal to the focused radius of the pulse at $z=0$, then the Raylegh diffraction limit is given by $Z_{R}=(\omega_{max}r_{a}^{2})/(2c)$, where $\omega_{max}=4c/a_{1}$ is the maximum effective angular frequency in the case of the fractional pulse.
+It is clear, then, that for a small value of the free parameter $a_{1}$ and a large aperture the region of abnormal (slow) decay can be very large.
+Generally, the larger the aperture, the longer the distance over which abnormally slow decay occurs.
+In a sense, this is analogous to an apertured Bessel beam, where the distance of non-diffracting propagation is not infinite but still increases with the size of the aperture.
+We can conclude that wave regions situated far from the propagation axis at $z=0$ contribute to the phenomenon of abnormally slow decay.
+As to the primitives in Eqs.
+(4) and (10), the electromagnetic and acoustic physical fields derived from them possess normal asymptotics.
+Therefore, these primitives are mainly of theoretical interest.
+6 Conclusion
+The solutions to the scalar wave equation with unusual asymptotic behavior in the far zone considered in this article are clearly of interest in mathematical physics, optics, electromagnetics and acoustics.
 It is remarkable that there are several solutions that decay with propagation distance more slowly than the common inverse proportional law, yet still possess finite energy.
@@
 Hence, the fields with abnormally slow decay in the far zone can in principle be implemented as near-cycle pulses of radiofrequency, optical, or acoustical waves.
-Figure 1: The real and imaginary parts of $U\left(\rho,z,t\right)$ at the instants $ct=0$ and $ct=4$ ($c\equiv 1$).
-The axis $x$ represents any axis transverse to the propagation axis $z$.
-The sign of the imaginary part has been reversed.
-The pulse width parameters are chosen as follows: $ct_{s}=0.3$ and $z_{s}=0.1$.
-Figure 2: ”Top view” of the 3D plot of the real part of $U\left(\rho,z,t\right)$ at four instants $t$. See caption of Fig. 1.
-Figure 3: The real and imaginary parts of $f\left(\rho,z,t\right)$ at $ct=0$ and $ct=10$.
-At positive times, the plot has mirror symmetry $z\rightarrow-z$ The pulse width parameters are $a_{1}=1$ and $a_{2}=2$.
-See also caption of Fig.
-1.
-Figure 4: Radial dependence of the imaginary parts of the expanding and converging splash waves defined in Eq.
-(21) and Eq.
-(22) at instants $ct=30$, $ct=170$, and $ct=-100$.
-The first two curves have been multiplied by $-R$, the third curve by $-50R$, and the last curve by $-R$.
-The pulsewidth parameter $ct_{s}=2$.
-Figure 5: Radial dependencies of the real parts of the waves defined in Eq.
-(24), Eq.
-(25), and Eq.
-(23) at instants $ct=30$ and $ct=160$.
-All curves have been multiplied by $R$.
-The pulse width parameter is $ct_{s}=2$.
-For comparison, a logarithmically diverging dependence is shown by the dotted curve, whereas the factor $1/2$ has been inserted in accordance with the presence of the same factor in Eqs.
-(24) and (25)
-Figure 6: Solid curves—time dependencies of the imaginary parts (multiplied by $ct$) of the fractional splash pulse $f\left(\rho=0,z,t\right)$ at points $z=30$, $z=90$, and $z=150$; dashed curve—an analogue of the expression under $lim$ in Eq.
-(2), i.e., $ct\cdot\Im f\left(\rho=0,z=ct+\Delta,t\right)$.
-For comparison, the dotted curve shows a diverging dependence $\sin(3\pi/8)\sqrt[4]{2ct}/2a_{1}$.
-The pulse width parameters are $a_{1}=3$ and $a_{2}=2$.
-$\Delta=-3/4$.
-The two last curves are shown only for the region $ct>0$.
-Figure 7: The $x$-component of the real electric field $E_{x}\left(x,y=0,z,t\right)$ at $ct=10$ calculated from the Hertz vector (1,1,0)$f\left(\rho,z,t\right)$.
-The pulse width parameters are $a_{1}=1$ and $a_{2}=2$.
-Cf., also, with Fig.
-3.
+The solutions presented here complement the extensive body of work on localized, nondiffracting, spatiotemporal, and autofocusing waves, which have already found applications in fields such as particle manipulation, micromachining, nonlinear spectroscopy, data communication and storage, microscopy, etc., up to medical diagnostics and therapy.
+Since, e.g., the fractional splash pulses share some essential properties with the listed waves and additionally exhibit exceptionally slow intensity decay and angular broadening during propagation, they have promising prospects for applications in the same fields.
+To summarize, we hope that abnormally decaying pulses as an emerging subfield in the study and applications of the localized space-time wave packets, which so far is represented to our best knowledge only by Refs.
+[3, 12, 5, 13, 14] and the present paper, will attract growing interest.
+APPENDIX.
+Proof of square integrability and finiteness of total energy of fractional splash pulses
+Here we present a thorough proof of the square integrability of fractional splash pulses and finiteness of their total energy.
+We rely on the approach developed in Refs.[7, 32], but our derivation of the key integral relation is simpler.
+A general splash pulse is given by a spectral superposition
+$f_{G}(\rho,z,t)=\int_{0}^{\infty}G_{FWM}(\rho,z,t)F_{\nu}(k)dk\;,$ (35)
+where
+$G_{FWM}(\rho,z,t)=\frac{\exp\left\{-k\rho^{2}/\left(a_{1}+i\tau\right)+ik\left(z+ct\right)\right\}}{a_{1}+i\tau}$ (36)
+is the wave function of the focus wave mode, $\tau=z-ct$, and $a_{1}$ is a positive constant with the dimension of lenght. $F_{\nu}(k)$ is a spectrum of the form
+$F_{\nu}(k)=\frac{1}{\Gamma\left(\nu+1\right)}k^{\nu}\exp\left(-a_{2}k\right);\;a_{2},k>0~,$ (37)
+which if $\nu=-1/4$ gives the 3/4 fractional splash pulse in Eq. (14).
+Following the approach of Refs.[7, 32], the square integrability of the function $f_{G}(\rho,z,t)$ is determined by finiteness of the right-hand side of the equality
+$\int\limits_{0}^{\infty}d\rho\rho\int\limits_{-\infty}^{\infty}dz~\left|f_{G}(\rho,z,t)\right|^{2}=\pi\int\limits_{0}^{\infty}dk~e^{2a_{1}k}E_{1}\left(2a_{1}k\right)\left|F_{\nu}(k)\right|^{2},$ (38)
+where $E_{1}\left(x\right)$ is the exponential integral function of the first order.
+For brevity we have omitted here the factor $2\pi$ which comes from integration over the azimuthal angle.
+So our task is to prove Eq.
+(38) and then to study whether the right-hand side is finite or diverges depending on the value of $\nu$.
+Inserting Eq.s (35)-(37) into left-hand side of Eq.
+(38), the integration over $\rho$ with setting $t=0$ (total energy does not depend on time), leaves the integration over $z$ into the form
+$I=\frac{1}{2}\int\limits_{-\infty}^{\infty}dz~\frac{\exp\left[-i\left(k_{1}-k_{2}\right)z\right]}{a_{1}\left(k_{1}+k_{2}\right)+i\left(k_{1}-k_{2}\right)z}~.$ (39)
+We evaluate this key integral by the following transformations.
+First for the sake of brevity we denote $\Delta k\equiv k_{1}-k_{2}$ and $\Lambda\equiv a_{1}\left(k_{1}+k_{2}\right)$.
+One can verify that $I=0$ unless $\Delta k=0$.
+But if we divide the integral into two parts, we obtain
+$I_{1}$ $\equiv\frac{1}{2}\int\limits_{-\infty}^{0}dz~\frac{\exp\left[-i\Delta kz\right]}{\Lambda+i\Delta kz}=\frac{i}{2}\frac{\operatorname{E}_{1}\left(\Lambda\right)e^{\Lambda}}{\Delta k},$ (40)
+$I_{2}$ $\equiv\frac{1}{2}\int\limits_{0}^{\infty}dz~\frac{\exp\left[-i\Delta kz\right]}{\Lambda+i\Delta kz}=-\frac{i}{2}\frac{\operatorname{E}_{1}\left(\Lambda\right)e^{\Lambda}}{\Delta k}.$ (41)
+These equalities can be found by packages of scientific calculations, or proved by changes of variables of integration.
+For example, Eq.
+(41) can be proved by the chain of the following changes of variables: $z\rightarrow\Lambda t,$ $t\rightarrow-it,$ $i(1+\Delta kt)\rightarrow y,$ $y\rightarrow it$, which results in
+$\frac{i}{2}\frac{e^{\Lambda}}{\Delta k}\int\limits_{1}^{\infty}\frac{\exp\left(-\Lambda t\right)}{t}=-\frac{i}{2}\frac{\operatorname{Ei}_{1}\left(\Lambda\right)e^{\Lambda}}{\Delta k},~Q.E.D.$ (42)
+Here, the definition of the exponential integral function $E_{1}\left(x\right)$ has been used. Eq. (40) is proved in the same way.
+Now we apply the relation
+$\lim_{\varepsilon\rightarrow 0}\frac{1}{x\mp i\varepsilon}=\pm i\pi\delta(x)+P\frac{1}{x}$ (43)
+known for distributions, to the right-hand sides of Eq.s (40) and (41).
+Common notations are used here: $\delta(x)$ is the Dirac delta and $P$ means the principal value.
+Applying Eq.
+(42) with the opposite signs to Eq.s (40) and (41) and summing them up, the principal values cancel out.
+Finally, returning to the initial designations we obtain
+$I=I_{1}+I_{2}=\pi\delta(k_{1}-k_{2})\operatorname{E}_{1}\left[a_{1}\left(k_{1}+k_{2}\right)\right]e^{a_{1}\left(k_{1}+k_{2}\right)}~.$ (44)
+From insertion Eq.s (35)-(37) into left-hand side of Eq.
+(38) there remains only the integration over $k_{1}$ and $k_{2}$, one of which can be trivially carried out thanks to the Dirac delta in the integrand.
+Denoting the other spectral variable by $k$, we obtain, finally, the right-hand side of Eq.
+(38).
+In order to study its finiteness, we make use of the upper bound $\exp(x)\operatorname{E}_{1}(x)<\ln(1+1/x)$ (Eq.
+5.1.20 in Ref.
+[33]).
+Hence, with the help of Eq.
+(37), a splash pulse of index $\nu$ is square integrable if the integral
+$B_{\nu}(a_{1},a_{2})\equiv\int\limits_{0}^{\infty}dk~\ln(1+1/2a_{1}k)\frac{\pi}{\Gamma^{2}\left(\nu+1\right)}k^{2\nu}\exp\left(-2a_{2}k\right)$ (45)
+is finite.
+With values $1/2<\nu\leq 0$ for the integral in Eq.
+(45) closed-form expressions can be found, which are more or less complicated combinations of elementary and special functions that have finite real values if $a_{1},a_{2}$ are positive parameters.
+If $v\leq-1/2,B_{\nu}(a_{1},a_{2})$ diverges, i.e., the 1/2-fractional pulse is not square integrable, although it exhibits abnormally slow decay with distance.
+Along these lines, also, finiteness of the scalar total energy of the fractional pulses with $1/2<\nu<0$ and of electromagnetic and acoustic fields derived from them can be established.
+However, since these energies contain spatial and temporal derivatives of the wave function $f_{G}(\rho,z,t),$ they are definitely finite without the need to calculate them, because derivatives mean multiplication by $k$ in the spectral domain, and thus remove the singularity at $k=0$ in Eq.
+(45) making the finiteness of it obvious.
+Square integrability and finiteness of the total energies, as well as the independence of them on time, have also been proved by direct numerical 2D integrations.
+For example, numerical integration in the left-hand side of Eq.
+(38), multiplied by $2\pi$ (due to the angular integral), in the case of a particular 3/4-fractional pulse ($\nu=-1/4,~a_{1}=1,~a_{2}=2$) yields a numerical result 29.041 independently from time instants $t=0,10,100$.
+For this case, the Mathematica package allows to find a closed-form expression for the right-hand side of Eq.
+(38), whose numerical value turns out to be equal to 29.041 as well.
+For the same case, the numerical value of the complicated closed-form expression of the integral in Eq.
+(45), multiplied also by $2\pi$, turns out to be 33.206, i.e., slightly larger as it should be for an upper bound.
+For the sake of completeness, we checked the finiteness of the total energy of the scalar field $f\left(\rho,z,t\right)$ also analytically.
+For this the modulus squared in the left-hand side of Eq.
+(38) was replaced by the scalar energy density
+$w=\frac{1}{2}\left|\frac{\partial}{\partial ct}f\left(\rho,z,t\right)\right|^{2}+\frac{1}{2}\left|\nabla f\left(\rho,z,t\right)\right|^{2}$ (46)
+and the right-hand side by the same quantity expressed through the spectral representation Eqs.
+(35)-(36).
+In the case of $\nu=-1/4$ the spectral integration in the righ-hand side results in a combination of $\sinh^{-1}$ and gamma functions, the numerical value of which for specified values of parameters ($a_{1}=1,~a_{2}=2$) turns out to be equal to the result of numerical integration on the left-hand side, i.e., which proves the finite value of the total energy.
+Finiteness of the total energy of the electromagnetic fields, derived from $f\left(\rho,z,t\right)$ by the Hertz vector technique, was also proved analytically with the help of Eq.
+(B14) in [32].
+Again, this conclusion is natural because the EM field expressions contain double derivatives of $f\left(\rho,z,t\right)$.
+[…]
+Funding No funding was used for the presented research.
+[…]
+Disclosures The authors declare no conflicts of interest.
+[…]
+Data availability No experimental data were generated or analyzed in the presented research.